# GPTKB: Building Very Large Knowledge Bases from Language Models

### v2

@@ Appendix A Further KB Content
Tables 3 and 4 provide further examples of GPTKB content, highlighting strengths as well as weaknesses. For Albuquerque, most triples appear correct, but 2 of the 3 sister cities are not, nor can several of the notablePerson values be verified. In the second case, most triples are correct, but the spouse is made up. In the online OpenReview appendix to this submission, we provide a larger sample of 10K triples.
Relation Object
annualEvent Albuquerque International Balloon Fiesta
area 189.5 square miles
areaCode 505
culturalEvent Albuquerque Film & Music Experience
New Mexico State Fair
Route 66 Summerfest
Albuquerque Pride
Gathering of Nations
elevation 5,312 feet
famousFor New Mexican cuisine
Film industry
Hot air ballooning
founded 1706
governmentType Strong mayor-council
instanceOf City
landmark ABQ BioPark
Old Town Albuquerque
Indian Pueblo Cultural Center
Sandia Peak Tramway
Petroglyph National Monument
locatedIn New Mexico
majorAirport Albuquerque International Sunport
nickname Duke City
notablePerson Larry David
Val Kilmer
Demi Moore
Jessica Chastain
Kurt Russell
Neil Patrick Harris
Billy the Kid
Vince Gilligan
Conan O’Brien
Georgia O’Keeffe
Walter White (fictional character)
Rudy Giuliani
Breaking Bad cast
population 560,000
river Rio Grande
sisterCity Torreón, Mexico
Chihuahua, Mexico
Hiroshima, Japan
sportsTeam Albuquerque Isotopes
New Mexico United
timeZone Mountain Standard Time
university University of New Mexico
zipCode 87101
@@
Table 3: Excerpt of GPTKB triples for subject Albuquerque.
Predicate Object
almaMater Georgia Institute of Technology
birthPlace Panama
children 2
genre Humor
instanceOf Person
knownFor PHD Comics
nationality American
occupation Cartoonist
Filmmaker
Writer
spouse Tanya K. K. Cham
website http://www.phdcomics.com
Table 4: Excerpt of GPTKB triples for subject Jorge Cham of PhDComics.

### v3

@@ Appendix A Further KB Content
Figures 5 and 6 provide further examples of GPTKB content. In the first case, all triples appear correct, but the predicate famousFor dominates the entity. In second case, most triples are correct, but the spouse is made up. In the online OpenReview appendix to this submission, we provide a larger sample of 10K triples.
Figure 5: Excerpt of GPTKB triples for subject Vienna.
@@
Figure 6: Excerpt of GPTKB triples for subject Jorge Cham of PhDComics.
Appendix B Class distribution by BFS-level
Figure 7 shows the most frequent classes by BFS-level. As one can see, diversity increases towards lower BFS-levels, and some classes, like awards, only emerge later.
Figure 7: Most frequent classes by BFS-level. The root level (#1) is excluded from the chart.
Appendix C Prompts
All prompts below are aided by OpenAI’s structured output feature, which allows defining a specific JSON schema, that the output has to adhere to. Therefore, the description of the output format in the prompts is less expansive than what is common in many other works.
Figure 8 contains the prompt used for triple elicitation. Figure 9 shows the prompt for NER. Figure 10 lists the prompt for taxonomy construction.
Figure 8: Prompt for knowledge elicitation.
You are an expert on named entity recognition (NER). Your task is to classify if given phrases are named entities (e.g., persons, organizations, works of art), or not (e.g., literals, dates, URLs, verbose phrases). Each phrase is given to you in a line.
Figure 9: Prompt for named-entity recognition (NER).
You are a knowledge base construction expert. Your task is to initialize a seed taxonomy with general categories, which you will update later with given classes. Please return only the seed taxonomy in json form with indentation. Class: <new class> You are a knowledge base construction expert. Your task is to create a taxonomy for a knowledge base. Beforehand, you need to give each given class a score describing how general it is. The score is an integer ranging only from 1, for the most general concept, to 10, for the most specific concept. Please return only the score of the given class. Candidate branches: <branches of current node> Class: <class to add> You are a knowledge base construction expert. Your task is to integrate a given class into the taxonomy. If the given class is a subclass of one of the candidate branches, return only the exact name of that branch. Otherwise, return only NULL. Taxonomy: <taxonomy from current node> Class: <class to add> You are a knowledge base construction expert. Your task is to update the given taxonomy with the given class. You can consider the categorization of the taxonomy, but you can not modify the names of the classes in the taxonomy. Please return only the updated taxonomy in JSON form.
Figure 10: Prompts for seed taxonomy construction (top), class generality scoring (upper middle), recursive class insertion check (lower middle), and (sub)taxonomy update (bottom).
Appendix D Analysis of temporal cutoff
In Figure 11, we plot the frequency of values for the year property. One can observe two things: 1) A steady increase of frequency with recency, likely mirroring the growth of web corpora over the years, which made more recent content overrepresented. 2) A sudden dropoff between 2023 (549 triples) and 2024 (75 triples), matching the self-declared knowledge cutoff of the GPT-4o-mini LLM.
1970197019701970198019801980198019901990199019902000200020002000201020102010201020202020202020202030203020302030002502502502505005005005007507507507501,00010001{,}0001 ,  [^0] YearCount
Figure 11: Number of triples for the year property, per year. The sudden drop from 2023 to 2024 matches the self-declared 2023 knowledge cutoff of GPT-4o-mini.
Appendix E Comparison of LLMs
In Table 4, we compare GPT-4o-mini with GPT-4o and Llama 3.1 70B in terms of accuracy. All LLMs are evaluated on a similar-sized set of entities (20K). One can observe significant accuracy differences, consistent with the order of their (estimated) parameter size, and general benchmark results.
GPT Llama GPT
4o-mini* 3.1 70B 4o
Web-verified triples 0.38 0.69 0.78
Entities on Wikidata 0.78 0.83 0.88
Web-verified entities 0.80 0.95 0.98
Table 4: Entity and triple verifiability comparison between different LLMs on a similar-sized set of entities. GPT-4o-mini* corresponds to the first 5 layers of what is reported in Figure 3.
Appendix F Sources for KB comparison
Sources of KB sizes in the comparison presented in Table 2: Wikidata: See https://www.wikidata.org/wiki/Wikidata:Statistics and https://grafana.wikimedia.org/d/000000175/wikidata-datamodel-statements. Wikidata5m: Wang et al. (2021). Yago4.5: Suchanek et al. (2024). DBpedia: English version as per Table 2 in Lehmann et al. (2015). NELL: As per Figure 5 (left) in Mitchell et al. (2018). Reverb: As per Lin et al. (2012) and https://web.archive.org/web/20220307185343/https://openie.allenai.org/.